A totl charge Q is broken in two parths Q1 and Q2 and they are placed at a distance R from each other the maximum force of repulsion between them will occur when

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

 let parts are sepereted at a distance of d

Q=q1+q2 (conservation of charge)

so repulsion force F= k Q1(Q2)/d2

                                      k Q1(Q-Q1)/d2

     here Q and d are constant so for maximum force differentiate the force w.r.t Q1

 dF/dQ1 = k/d2 [Q -2Q1]

 for maximum force dF/dQ1=0

                             Q-2Q1=0

                              Q1 +Q2 -2Q1 =0

                             Q1 =Q2

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin